# LATTICES OF FLATS FOR SYMPLECTIC MATROIDS

### v5

@@ 4.10. Theorem:
Let $\mathcal{L}$ be a $C_{n}$ lattice and $\mathcal{P}$ the induced geometric lattice then:
$\mathcal{L}=\mathcal{P}\cap P^{ad}(J)$
@@
Assume $X$ is an admissible element of $\mathcal{P}$ that covers (in $\mathcal{P}$) an element $Y$ in $\mathcal{L}$. As the union of the elements of $\mathcal{L}$ covering $Y$ in $\mathcal{L}$ is $J\setminus(Y\cup Y^{*})$ we must have some element $Z$ in $\mathcal{L}$ covering $Y$ and $X\cap Z\in\mathcal{L}$ covering $Y$. Therefore, $Y\subsetneq X\subsetneq Z$ and we have changed the rank of $X$ which contradicts theorem [3.8]. As $A(\mathcal{L})=A(\mathcal{P})$, if we had $\mathcal{L}\neq\mathcal{P}\cap P^{ad}(J)$, there would have been such $X$.
We can now use theorems [4.8], [4.9], and [4.10] to formulate the characterization of symplectic matroids corresponding to $C_{n}$ lattices using rank functions.
@@ 4.12. Theorem:
Let $\mathcal{I}$ be a family of independent sets of a symplectic matroid of rank $d$ with no loops, then $\mathcal{I}$ corresponds to a $C_{n}$ lattice and denoted a ranked symplectic matroid iff the function $r:P\left(J\right)\rightarrow\mathbb{Z}_{+}$, sending a set $A$ to $min(d,\rho(A))$, is the rank function of a loop-free ordinary matroid on the ground set $J$ for the following function $\rho$:
@@
The "only if" direction is Theorem [4.8]. For the "if" direction it is enough to show, again using theorem [4.8], that $\mathcal{P}\cap P^{ad}(J)$ is a $C_{n}$ lattice for the induced geometric lattice $\mathcal{P}$ of $r$. We note that every atom of $\mathcal{P}$ is admissible by the definition of $r$ and $\mathcal{I}$ being without loops. Therefore, we are reduced to showing that any non-maximal admissible flat $F$ of $\mathcal{P}$ is covered by exactly one inadmissible flat $F\cup F^{*}$. To see that any inadmissible flat covering $F$ is contained in $F\cup F^{*}$, let $I\cup\left\{a\right\}$ be an independent set spanning $F$ such that $I\cup\left\{a^{*}\right\}$ is also independent. If $b,b^{*}\notin F$, then we must have $b,b^{*}\notin cl_{r}(I\cup\left\{a,a^{*}\right\})$. As both $I\cup\left\{a,b\right\}$ and $I\cup\left\{a,b^{*}\right\}$ are independent and $I\cup\left\{a\right\}\in\mathcal{I}^{\triangle}$, contradicting $cl_{r}(I\cup\left\{a,a^{*}\right\}),cl_{r}(I\cup\left\{b,b^{*}\right\})$ being of the same rank. As the rank function $r$ does not allow independent sets with more than one pair of inadmissible elements, we must also have $F\cup F^{*}$ contained in the inadmissible flat covering $F$.
Proof of theorem 4.5:
By definition, $NBB$ is a hereditary property, so $I\left(\mathcal{L},P\right)$ is a subset closed family.
  1. (1)
    Given a transversal $T\subset J$ we have by lemma [4.5] that $I\left(\mathcal{L},P\right)\cap T=I\left(\mathcal{L}\cap T,P\mid_{T}\right)$ and the restriction of $P$ to $T$ the set of all linear orders on $T$. By Theorem [3.6] we then have $I\left(\mathcal{L},P\right)\cap T$ the family of independent sets of an ordinary matroid with a ground set $T$.
  2. (2)
    Let $I_{1},I_{2}\in I^{ad}\left(\mathcal{L},P\right)$ be such that $\left|I_{1}\right|\lneq\left|I_{2}\right|$. As $I_{1},I_{2}\in I\left(\mathcal{L},P\right)$, which is a family of independent sets of a matroid, there exist $a\in I_{2}$ such that $I_{1}\cup\left\{a\right\}\in I\left(\mathcal{L},P\right)$. Assume all such $a$ has $a^{*}\in I_{1}$ and observe that $\left|I_{1}\setminus I_{2}^{*}\right|\lneq\left|I_{2}\setminus I_{1}^{*}\right|$ for the smaller independent sets $I_{1}\setminus I_{2}^{*},I_{2}\setminus I_{1}^{*}$. Let $b\in I_{2}\setminus I_{1}^{*}$ be such that $(I_{1}\setminus I_{2}^{*})\cup\left\{b\right\}\in I^{ad}\left(\mathcal{L},P\right)$. We actually have $I_{1}\setminus I_{2}^{*}\in I\left(\mathcal{L},P\right)$ because we have removed all inadmissible pairs in $I_{1}\cup I_{2}$. Using Lemma [4.7], we see that, as $I_{2}\cup\left\{a\right\}\in I\left(\mathcal{L},P\right)$ is inadmissible, $rank(I_{2})\lneq rank(\mathcal{L})$. Using Lemma [4.7] once more, we get $I_{1}\cup\left\{b^{*}\right\}\in I\left(\mathcal{L},P\right)$ satisfying the second part of condition (2) of definition [4.3].
Using ordinary matroid terminology, we have defined a "cryptomorphism" $f$ sending a $C_{n}$ lattice to the symplectic matroid defined by its family of admissible independent sets. By Theorems [3.7] and [4.10] the family of admissible independent sets is uniquely determined by the induced geometric lattice, and so we have $f$ injective, see remark [4.1].
@@ 4.13. Remark 1:
An ordinary matroid on $J$ of rank $\geq 3$ with admissible bases, and so also a symplectic matroid, is never a ranked symplectic matroid. The rank function from Theorem [4.11] will never be submodular.
@@ 4.14. Remark 2:
If $\mathcal{B}$ is an ordinary matroid with a partition of its atoms into pairs $1,1^{*},...,k,k^{*}$ such that its rank function identifies with the induced rank function of its admissible independent sets (in the sense of Theorem [4.11]) then $\mathcal{B}\cap P^{ad}(J)$ is a ranked symplectic matroid.
The second remark introduces a new way of constructing families of symplectic matroids using ordinary matroids. This can be helpful because the theory of ordinary matroids is much more developed. We show an example of how any spike with no tip gives rise to a ranked symplectic matroid.
@@ 4.15. Definition:
The cycles of $\mathcal{C}$ are called balanced and the other cycles are unbalanced. The lift matroid represented by $\mathcal{M}(G,\mathcal{C})$ has for its ground set the set of edges of $G$ and, as independent sets, the sets of edges containing at most one cycle, which is unbalanced. A spike with no tip is a lift matroid with $G$ the graph obtained from the cycle on $n$ vertices by doubling every edge and $\mathcal{C}$ some set of cycles that do not contain any pair of double edges satisfying the theta property.
To see that any spike with no tip $\mathcal{M}(G,\mathcal{C})$ is a simple matroid of the form defined in Theorem [4.8], notice that $J=[n]\sqcup[n^{*}]$ with $n$ being the size of the original cycle in $G$. Every simple cycle in $G$ is a set of the form $\left\{i,i^{*}\right\}$ for some edge $i$ or a transversal $T$. Therefore, every non maximal admissible independent set $I$ of $\mathcal{M}(G,\mathcal{C})$ does not contain a cycle and $I\cup\left\{i^{*}\right\}$ is an independent set of $\mathcal{M}(G,\mathcal{C})$ for every $i\in I$. Moreover, as $\left\{i,i^{*},j,j^{*}\right\}$ contains two cycles it cannot be extended to an independent set.
$v$$u$$1^{*}$$2^{*}$$\{1,2,1^{*},2^{*}$}$\{1,2^{*}\}$$\{1^{*},2\}$$\emptyset$$\{1,2\}$$\{1^{*},2^{*}\}$($a$)($b$)($c$)
@@ 4.16. Figure:
In [4] T. Chow introduced graphic symplectic matroids. We note that our construction of spikes with no tip is not necessarily graphic symplectic matroids, as the transversal being even does not correspond to the number of starred elements.
@@ 5.1. Definition:
A simplicial complex $\Delta$ is a set of simplices satisfying the following conditions:
@@ 5.2. Definition:
Let $\Delta$ be a finite simplicial complex. We say that $\Delta$ is pure $d$ dimensional if all its facets are of dimension $d$. A pure $d$ dimensional simplicial complex $\Delta$ is said to be shellable if its facets can be ordered $F_{1},...,F_{t}$ in such a way that:
@@
Is a pure $\left(d-1\right)$ dimensional complex for $j=2,3,...,t$, with:
@@
To a finite poset $\mathcal{L}$ one can associate a simplicial complex $\Delta\left(\mathcal{L}\right)$ (the order complex) of all unrefinable chains of $\mathcal{L}$. Also, if $\mathcal{L}$ is a graded poset of rank $d$ then $\Delta\left(\mathcal{L}\right)$ is pure $d$ dimensional. We say that a finite and graded poset $\mathcal{L}$ is shellable if its order complex $\Delta\left(\mathcal{L}\right)$ is shellable.
@@ 5.3. Definition:
    For all $j=2,...,t$ we have $\left[A_{j},\hat{1}\right]$ admit a recursive atom ordering in which the atoms of $\left[A_{j},\hat{1}\right]$ that come first in the ordering are those that cover some $A_{i}$ where $i<j$.
@@ 5.4. Theorem:
In [7] A. Bjorner and M. Wachs proved that if a graded poset $L$ admits a recursive atom ordering, then the order complex $\Delta\left(L\right)$ is shellable.
@@ 5.5. Remark:
It is possible to prove a stronger type of lexicographic shellability called EL-shellability by taking the known EL-labeling for geometric lattices (can be found in [10]) of its enveloping ordinary matroid, starting with the atoms of an admissible coatom. We decided to prove CL-shellability, equivalent to admitting a recursive atom ordering, because we obtain more shelling orders and the characterization in part 6.
@@ 5.6. Lemma:
Assume that $A,B$ are atoms of $\mathcal{L}$ with $B^{*}\cap A\neq\emptyset$. If $B^{*}\setminus A\neq\emptyset$ we must have for every $c\in A\setminus B^{*}$ some element $c\in C\in\mathcal{L}$ with $C$ covering $B$, but as we proved that $\mathcal{L}$ is atomic and so we must have $A\subseteq C$ which is a contradiction as:
$B^{*}\cap A\neq\emptyset\Rightarrow C^{*}\cap A\neq\emptyset$
$C$ will not be admissible and not of full rank as $\mathcal{L}$ is of rank $\geq 3$. So we must have:
$B^{*}\cap A\neq\emptyset\Rightarrow B^{*}=A$
And so if $a\in A$ there exist some atom $B\in\mathcal{L}$ such that $a^{*}\in B$ (as the set of atoms partition $J$), but then $B=A^{*}$.
@@ 5.7. Lemma :
Let $\mathcal{L}$ be a $C_{n}$ lattice and $A\in\mathcal{L}$ then the restriction $\left[A,\hat{1}\right]\subseteq\mathcal{L}$ is order isomorphic to a $C_{n}$ lattice on the ground set $J^{\prime}=J\setminus\left(A\cup A^{*}\right)$.
@@
We define
$\left[A,\hat{1}\right]\setminus A=\left(\left\{X\setminus A\mid X\in\left[A,\hat{1}\right]\right\}\setminus\left\{J\setminus A\right\}\right)\cup\left\{J^{\prime}\right\}$
This is obviously isomorphic to $\left[A,\hat{1}\right]$ by the lattice morphism:
$\psi:\left[A,\hat{1}\right]\rightarrow\left[A,\hat{1}\right]\setminus A$
$X\mapsto X\setminus A$
(At first glance, it is not clear that $\left[A,\hat{1}\right]\setminus A$ and $\left[A,\hat{1}\right]$ are lattices, but this becomes obvious by proving $\left[A,\hat{1}\right]\setminus A$ is a $C_{n}$ lattice).
We continue with proving the four conditions:
@@
    Every element $B\in\left[A,\hat{1}\right]\setminus A$ is admissible as it is an element of $\mathcal{L}$ with some elements removed.
@@
    $\forall B,C\in\left[A,\hat{1}\right]\setminus A$ we have:
    $A\subseteq\left(A\cup B\right)\cap\left(A\cup C\right)\in\mathcal{L}\Rightarrow\left(A\cup B\right)\cap\left(A\cup C\right)\in\left[A,\hat{1}\right]\Rightarrow A\cap B\in\left[A,\hat{1}\right]\setminus A$
@@
    For $B\in\left[A,\hat{1}\right]\setminus A$ let $\left\{B_{1},...,B_{m}\right\}$ the set of elements in $\left[A,\hat{1}\right]\setminus A$ covering $B$. Therefore, we have $B\cup A\in\mathcal{L}$ and a set of elements $\left\{C_{1},...,C_{k}\right\}$ covering $B\cup A$ in $\mathcal{L}$. To see $C_{i}\setminus A$ covers $B$ in $\left[A,\hat{1}\right]\setminus A$, assume that there is some $C\in\left[A,\hat{1}\right]\setminus A$ such that $B\leq C\leq C_{i}\setminus A$. hence $C\cup A\in\mathcal{L}$ and so $B\cup A\leq C\cup A\leq C_{i}$ in $\mathcal{L}$. Resulting in $C\cup A=B\cup A\Rightarrow C=B$ or $C\cup A=C_{i}\Rightarrow C=C_{i}\setminus A$. In a similar way, we show that $B_{i}\cup A\in\mathcal{L}$ covers $B\cup A$ in $\mathcal{L}$ and so we get:
    $\left\{B_{1},...,B_{m}\right\}=\left\{C_{1}\setminus A,...,C_{m}\setminus A\right\}$
    Therefore, we have:
    $B_{i}\cap B_{j}=\left(C_{i}\setminus A\right)\cap\left(C_{j}\setminus A\right)=\left(C_{i}\cap C_{j}\right)\setminus A=\left(B\cup A\right)\setminus A=B$
@@
    $\cup_{i=1}^{m}B_{i}=\cup_{i=1}^{m}\left(C_{i}\setminus A\right)=\left(\cup_{i=1}^{m}C_{i}\right)\setminus A=\left(J\setminus\left(B\cup A\right)^{*}\right)\setminus A=J^{\prime}\setminus B^{*}$

### v6

@@ 4.12. Theorem:
Let $\mathcal{L}$ be a $C_{n}$ lattice, and $\mathcal{P}$ denote the geometric lattice induced by $\mathcal{L}$. Then $\mathcal{P}$ contains no additional admissible elements beyond those already in $\mathcal{L}$; that is, if $X\in\mathcal{P}$ is admissible, then necessarily $X\in\mathcal{L}$.
@@
We first prove that if $I\in I^{ad}(\mathcal{L})\setminus I^{\overline{ad}}(\mathcal{L})$ then $\vee_{\mathcal{P}}I$ is inadmissible.
Using Lemma 3.3, let $a\in I$ be such that $I\setminus\left\{a\right\}\in I^{\overline{ad}}(\mathcal{L})$. As a result of property 4 of Definition 2.3 we must have $a^{*}\in\vee_{\mathcal{L}}(I\setminus\left\{a\right\})$. Using Lemma 4.10 we have
$rank_{\mathcal{L}}(I\setminus\left\{a\right\})=\left|I\setminus\left\{a\right\}\right|=rank_{\mathcal{P}}(I\setminus\left\{a\right\})\Rightarrow a^{*}\in\vee_{\mathcal{P}}(I\setminus\left\{a\right\})\subseteq X$
To prove the Theorem, let $X$ be an admissible element of $\mathcal{P}$, and let $I$ be an independent set such that $\vee_{\mathcal{P}}I=X$. We must then have $I\in I^{\overline{ad}}(\mathcal{L})$ and by Lemma 4.10, $rank_{\mathcal{L}}(I)=\left|I\right|=rank_{\mathcal{P}}(I)$. As $\mathcal{P}$ is closed under intersection, we have
$I\subseteq(\vee_{\mathcal{P}}I)\cap(\vee_{\mathcal{L}}I)\in\mathcal{P}\Rightarrow rank_{\mathcal{P}}((\vee_{\mathcal{P}}I)\cap(\vee_{\mathcal{L}}I))=rank_{\mathcal{P}}(X)\Rightarrow X=\vee_{\mathcal{L}}I\in\mathcal{L}$
We can now use Theorems 4.9, 4.11, and 4.12 to formulate the characterization of symplectic matroids corresponding to $C_{n}$ lattices using rank functions.
@@ 4.14. Theorem:
Let $\mathcal{I}$ be a family of independent sets of a symplectic matroid of rank $d$ with no loops, then $\mathcal{I}$ corresponds to a $C_{n}$ lattice and denoted a ranked symplectic matroid if and only if the function $r:P\left(J\right)\rightarrow\mathbb{Z}_{+}$ is the rank function of a loop-free ordinary matroid on the ground set $J$ for the following function $r$:
@@
The "only if" direction is exactly Theorem 4.9 and 4.11, as the rank of a set in an ordinary matroid is the size of a maximal independent set contained in it. For the "if" direction it is enough to show, again using Theorem 4.11, that the admissible part of $\mathcal{P}$ is a $C_{n}$ lattice for the induced geometric lattice $\mathcal{P}$ of $r$. We note that every atom of $\mathcal{P}$ is admissible by the definition of $r$ and $\mathcal{I}$ being without loops. Therefore, we are reduced to showing that any non-maximal admissible flat $F$ of $\mathcal{P}$ is covered by exactly one inadmissible flat $F\cup F^{*}$. To see that any inadmissible flat covering $F$ is contained in $F\cup F^{*}$, let $I\cup\left\{a\right\}$ be an independent set spanning $F$ such that $I\cup\left\{a^{*}\right\}$ is also independent. If $b,b^{*}\notin F$, then we must have $b,b^{*}\notin\vee_{r}(I\cup\left\{a,a^{*}\right\})$, with $\vee_{r}$ being the join in the geometric lattice corresponding to $r$. As both $I\cup\left\{a,b\right\}$ and $I\cup\left\{a,b^{*}\right\}$ are independent and $I\cup\left\{a\right\}\in\mathcal{I}^{\triangle}$, contradicting $\vee_{r}(I\cup\left\{a,a^{*}\right\}),\vee_{r}(I\cup\left\{b,b^{*}\right\})$ being of the same rank. As the rank function $r$ does not allow independent sets with more than one pair of inadmissible elements, we also must have $F\cup F^{*}$ contained in the inadmissible flat covering $F$.
Using ordinary matroid terminology, we have defined a "cryptomorphism" $f$ sending a $C_{n}$ lattice to the symplectic matroid defined by its family of admissible independent sets. By Theorems 3.7 and 4.9 the family of admissible independent sets is uniquely determined by the induced geometric lattice, and so we have $f$ injective; see Remark 4.1.
@@ 4.15. Remark 1:
An ordinary matroid on $J$ of rank $\geq 3$ with admissible bases, and so also a symplectic matroid, is never a ranked symplectic matroid. The rank function from Theorem 4.14 will never be submodular.
@@ 4.16. Remark 2:
In Theorem 4.14 we have required $\mathcal{I}$ to be a family of independent sets of a symplectic matroid, we can weaken this condition, the same proof will work for the following Theorem:
Let $\mathcal{M}$ be an ordinary matroid of rank $d$ on the ground set $J$. If $\mathcal{M}$ contains an admissible base and its rank function is the one described in Theorem 4.14 then $\mathcal{L}$, the admissible part of the geometric lattice $\mathcal{P}$ corresponding to $\mathcal{M}$, is a $C_{n}$ lattice. Moreover, the ranked symplectic matroid corresponding to $\mathcal{L}$ is the family of admissible bases of $\mathcal{M}$.
Remark 4.16 introduces a new way of constructing families of symplectic matroids using ordinary matroids. This can be helpful because the theory of ordinary matroids is much more developed. We show an example of how any spike with no tip gives rise to a ranked symplectic matroid.
@@ 4.17. Definition:
The cycles of $\mathcal{C}$ are called balanced and the other cycles are unbalanced. The lift matroid represented by $\mathcal{M}(G,\mathcal{C})$ has for its ground set the set of edges of $G$ and, as independent sets, the sets of edges containing at most one cycle, which is unbalanced. A spike with no tip is a lift matroid with $G$ the graph obtained from the cycle on $n$ vertices by doubling every edge and $\mathcal{C}$ some set of cycles that do not contain any pair of double edges and satisfying the theta property.
4.18. Theorem:
The family of admissible bases of any spike with no tip, $\mathcal{M}(G,\mathcal{C})$, constitutes a ranked symplectic matroid.
Proof:
We will show that $\mathcal{M}(G,\mathcal{C})$ is a simple matroid admitting the properties described in Remark 4.16.
First, notice that $J=[n]\sqcup[n^{*}]$ with $n$ being the size of the original cycle in $G$, labeling one of the edges $\left\{x,y\right\}$ with $i\in[n]$ gives the other $\left\{x,y\right\}$ edge the label i-i.
Every simple cycle in $G$ is a set of the form $\left\{i,i^{*}\right\}$ for some edge $i$ or a transversal $T$. Therefore, every maximal independent set of $\mathcal{M}(G,\mathcal{C})$ is an unbalanced transversal $T$ or a missing transversal of the form $(T\setminus\left\{i\right\})\cup\left\{j\right\}$, for a transversal $T$, and $i,j^{*}\in T$.
To see that there always exists an unbalanced transversal, we observe that, for example, $n],([n]\setminus\left\{1\right\})\cup\left\{1^{*}\right\}$ cannot be both balanced. Otherwise, we will have $\left\{1,1^{*}\right\}$ balanced by the Theta property. Thus, we have found an admissible basis for $\mathcal{M}(G,\mathcal{C})$.
We are left to show that the rank function of $\mathcal{M}(G,\mathcal{C})$ agrees with the rank function $r$ defined in Theorem 4.14. Recall that for an ordinary matroid, the rank of a set is the size of a maximal independent set contained in it. If $X\subseteq J$ is admissible, then trivially $rank_{\mathcal{M}(G,\mathcal{C})}(X)=r(X)$. This is also the case if $X$ contains an unbalanced transversal. Otherwise, let $I\subseteq X$ be a maximal admissible independent set, if $I^{*}\cap X\neq\emptyset$ then for $i\in I^{*}\cap X$ we have $I\cup\left\{i\right\}$ independent in $\mathcal{M}(G,\mathcal{C})$ and $rank_{\mathcal{M}(G,\mathcal{C})}(X)=\left|I\right|+1$. Every non-maximal admissible independent set of $\mathcal{M}(G,\mathcal{C})$ is delta-independent we have:
$rank_{\mathcal{M}(G,\mathcal{C})}(X)=\left|I\right|+1=\left|I\setminus\left\{i\right\}\right|+2=r(X)$
[…]
@@ 4.19. Figure:
In [4] T. Chow introduced graphic symplectic matroids. We note that our construction of spikes with no tip is not necessarily a graphic symplectic matroids, as the transversal being even does not correspond to the number of starred elements.
@@ 5.1. Definition:
A simplicial complex $\Delta$ is a set of simplices that satisfy the following conditions:
@@ 5.2. Definition:
Let $\Delta$ be a finite simplicial complex. We say that $\Delta$ is pure $d$ dimensional if all its facets (inclusion maximal faces) are of dimension $d$. A pure $d$ dimensional simplicial complex $\Delta$ is said to be shellable if its facets can be ordered $F_{1},...,F_{t}$ in such a way that:
@@
Is a pure $\left(d-1\right)$ dimensional complex for j=2,3,,tj=2,3,...,t, with:
@@
To a finite poset $\mathcal{L}$ one can associate a simplicial complex $\Delta\left(\mathcal{L}\right)$ (the order complex of $\mathcal{L}$) of all chains of $\mathcal{L}$. Also, if $\mathcal{L}$ is a graded poset of rank $d$, then $\Delta\left(\mathcal{L}\right)$ is pure $d$ dimensional. We say that a finite and graded poset $\mathcal{L}$ is shellable if its order complex $\Delta\left(\mathcal{L}\right)$ is shellable.
@@ 5.3. Definition:
    For all j=2,,tj=2,...,t we have $\left[A_{j},\hat{1}\right$] admit a recursive atom ordering in which the atoms of $\left[A_{j},\hat{1}\right$] that come first in the ordering are those that cover some $A_{i}$ where $i<j$.
@@ 5.4. Theorem:
Let $P$ be a graded poset. If $P$ admits a recursive atom ordering, then the order complex $\Delta\left(P\right)$ is shellable.
@@ 5.5. Remark:
It is possible to prove a stronger type of lexicographic shellability called EL-shellability by taking the known EL-labeling for geometric lattices (can be found in [10]) of its enveloping ordinary matroid, starting with the atoms of an admissible coatom. To expand the range of shelling orders and support the characterization in part 6, we chose to establish CL-shellability, which is equivalent to admitting a recursive atom ordering.
@@ 5.6. Lemma:
Assume that $A,B$ are atoms of $\mathcal{L}$ with $B^{*}\cap A\neq\emptyset$. Assume by negation that $A\setminus B^{*}\neq\emptyset$ and let $a\in A\setminus B^{*}$. Applying property $4$ of Definition 2.3 we have $a\notin B$ and an element $C\in\mathcal{L}$ containing $a$ and covering $B$. As $a\in C\cap A\in\mathcal{L}$, we have $C\cap A=A\Rightarrow A\subseteq C$. By our assumption, we now have:
$A\cap B^{*}\neq\emptyset\overset{B\subseteq C}{\Rightarrow}A\cap C^{*}\neq\emptyset\overset{A\subseteq C}{\Rightarrow}C\cap C^{*}\neq\emptyset$
That is, $C$ is inadmissible. As $rank(C)=2\lneq rank(\mathcal{L)}$ and $J$ is the only inadmissible element in $\mathcal{L}$, this is a contradiction. Therefore, $B^{*}\cap A\neq\emptyset\Rightarrow B^{*}=A$.
@@ 5.7. Lemma :
Let $\mathcal{L}$ be a $C_{n}$ lattice and $A\in\mathcal{L}$ then the restriction $\left[A,J\right]\subseteq\mathcal{L}$ is order isomorphic to a $C_{n}$ lattice on the ground set $J^{\prime}=J\setminus\left(A\cup A^{*}\right)$.
@@
Let $A\in\mathcal{L}$, we will work with the lattice morphism:
$\psi:\left[A,J\right]\rightarrow 2^{J^{\prime}}$
$X\mapsto\begin{cases}J^{\prime}&X=J\\ X\setminus A&X\neq J\end{cases}$
We observe that $\psi(X)\subseteq J^{\prime}$ for any $X\in\left[A,J\right]$ because $X\cap A^{*}=\emptyset$ or $X=J$. It is also clear that $\psi$ is injective, and hence it is a lattice isomorphism on its image.
We continue by proving the four properties of Definition :
@@
    Every element $B\in\psi(\left[A,J\right])$ except for $J^{\prime}$ is admissible as it is a subset of a proper element of $\mathcal{L}$.
@@
    $\forall B,C\in\psi(\left[A,J\right])$ we have:
    $A\subseteq\left(A\cup B\right)\cap\left(A\cup C\right)\in\left[A,J\right]\Rightarrow B\cap C\in\psi(\left[A,J\right])$
@@
    For $J^{\prime}\neq B\in\psi(\left[A,J\right])$ let $\left\{B_{1},...,B_{m}\right\}$ be the set of elements in $\psi(\left[A,J\right])$ covering $B$. If $\left\{B_{1},...,B_{m}\right\}=\emptyset$ or $\left\{B_{1},...,B_{m}\right\}=J^{\prime}$, then property $(4)$ follows directly. We now consider the nontrivial case where $|\left\{B_{1},...,B_{m}\right\}|\geq 2$. We know that $\psi$ is a lattice isomorphism from an upper interval, hence for any $X\in\left[A,J\right]$, we have $Y$ cover $X$ in $\mathcal{L}$ if and only if $\psi(Y)$ covers $\psi(X)$ in $\psi(\left[A,J\right])$. Therefore, we have $\psi^{-1}(B)=B\cup A\in\mathcal{L}$ and a set of elements $\left\{\psi(B_{1}),...,\psi(B_{m})\right\}=\left\{B_{1}\cup A,...,B_{m}\cup A\right\}$ covering $B\cup A$ in $\mathcal{L}$. As a result of property $(4)$ for $\left\{B_{1}\cup A,...,B_{m}\cup A\right\}$ in $\mathcal{L}$ we have:
    $B_{i}\cap B_{j}=((B_{i}\cup A)\cap(B_{j}\cup A))\setminus A=(B\cup A)\setminus A=B$
@@
    $\cup_{i=1}^{m}B_{i}=(\cup_{i=1}^{m}(B_{i}\cup A))\setminus A=(J\setminus A^{*})\setminus A=J^{\prime}$